OpenStudy (karatechopper):

For problems 1–5, simplify the expression. Answers written in decimal form will not be accepted. Each of these problems is worth 1 point. "v/" is a radical by the way ___ 1.v/96 ______ 2. 8 v/63x^5 ___________ 3. v/128x^5y^2 ___ 4. ^3v/32 ________ 5. ^3v/56x^14 @poopsiedoodle

OpenStudy (karatechopper):

HELP HIM----> @poopsiedoodle

OpenStudy (anonymous):

Break them down so that you have something inside the radical that you can simplify. For example: \[\sqrt{96} = \sqrt{16\times6} = \sqrt{16}\times \sqrt{6} = 4 \sqrt{6}\]

OpenStudy (poopsiedoodle):

I already have that one, but the ones with the exponents on the outside and the variables are confusing the poopy out of me.